fins the critical value ta/2 that corresponds to 99% confidence level for a sample size n=15
Answer #1

Solution :

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

n = 15

df = 15 - 1 = 14

t 0.005,14 = 2.977